OpenStudy (anonymous):

Which of the following could be used to calculate the area of the sector in the circle shown above? π(10in)^2 37 over 360 <---- my answer π(10in)37 over 360 π(37in)210 over 360 π(37in)10 over 360

OpenStudy (anonymous):

@mathmath333

OpenStudy (jdoe0001):

hmmm what lead you to think so?

OpenStudy (anonymous):

well i thought it was that because since r= 10 but i was a little unsure if it was right

OpenStudy (jdoe0001):

well \(\bf \textit{sector of a circle}=\cfrac{\theta\pi r^2}{360}\qquad \begin{cases} \theta=37\\ r=10 \end{cases}\implies \cfrac{37\pi 10^2}{360}\)

OpenStudy (anonymous):

oh ok so then it would be C right ?

OpenStudy (jdoe0001):

hmmm ahemm \(\bf \textit{sector of a circle}=\cfrac{\theta\pi r^2}{360}\qquad \begin{cases} \theta=37\\ r=10 \end{cases}\implies \cfrac{37\pi 10^2}{360} \iff \cfrac{\pi 10^2 37}{360}\) commutative property

Miracrown (miracrown):

You're right @Moo_Moo17 It is A indeed. If angle is 360 then we get the area of the whole circle. It should be proportional to the square of the radius. \[S = r ^{2}37\] And there should be angle of the sector and we know that for the circle the area is: \[\pi r^2\] So, if our sector is 360 degrees we have to get full circle.

OpenStudy (anonymous):

oh ok thank yo so much miracrown ^w^

Miracrown (miracrown):

yw :)